OpenStudy (anonymous):

Please help....(They wouldnt let me study this, so im counting on you to even pass :/)

OpenStudy (anonymous):

In the figure, j || k and m<1 = 63 degrees. What is m<4? ____ degrees.

OpenStudy (anonymous):

@SaltLife23 @M4thM1nd @YanaSidlinskiy

OpenStudy (anonymous):

you can see that m<1 + m<4 = 180 deg so... m<4 = 180 - 63 m<4 = 117 deg

OpenStudy (anonymous):

Oooooh.... Thanks! :)

OpenStudy (anonymous):

Np

OpenStudy (anonymous):

If you have more, I can help (:

OpenStudy (anonymous):

@Angel_Bear19 Alright, i have 2 more, ill ping you to them :)